OpenStudy (anonymous):

The maximum height of a vehicle that can safely pass under a bridge is 12 feet 5 inches. A truck measures 162 inches in height. Which best explains whether or not the truck can pass safely under the bridge?

OpenStudy (anonymous):

The truck can pass safely under the bridge. The truck is 13 inches shorter than the maximum height. The truck cannot pass safely under the bridge. The truck is 13 inches taller than the maximum height. The truck can pass safely under the bridge. The truck is 12 inches shorter than the maximum height. The truck cannot pass safely under the bridge. The truck is 12 inches taller than the maximum height

OpenStudy (anonymous):

there are 12 inches in a foot, so the height of the bridge is 12*12=144+5=149 162-149=13 so the answer is b.